archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    00 54AM Reason for closing Done Comments 7 Comment by OVH Tuesday 25 February 2014 00 39AM We will upload the new release on the secondary partition of the router Comment by OVH Tuesday 25 February 2014 00 40AM Erasing the alternate release Comment by OVH Tuesday 25 February 2014 00 40AM Release erase completed Installing la nouvelle release Comment by OVH Tuesday 25 February 2014 00 40AM Installation completed

    Original URL path: http://status.ovh.es/?do=details&id=6412 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Complete Details We are having an issue on the collect 57 We are opening a ticket with the operator Date Saturday 22 February 2014 22 04PM Reason for closing Done Comments 2 Comment by OVH Saturday 22 February 2014 22 02PM The ticket is opened The operator has been notified of the incident on the network Their team are investigating Comment by OVH Saturday 22 February 2014 22 03PM The

    Original URL path: http://status.ovh.es/?do=details&id=6395 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Core Status CLOSED Percent Complete Details Following the optic intervention http status ovh net do details id 6137 Canteleu CAN59 and Haubourdin HAU59 sites will be isolated during the intervention Date Thursday 13 February 2014 14 34PM Reason for closing

    Original URL path: http://status.ovh.es/?do=details&id=6320 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    at around 5 00am Only clients with the LNS Paris as a primary LNS but currently on the LNS backup are concerned More information http status ovh net do details id 6263 http status ovh net do details id 6205 Date Friday 14 February 2014 15 02PM Reason for closing Done Comments 2 Comment by OVH Wednesday 12 February 2014 12 39PM The sessions were successfully migrated We monitor the

    Original URL path: http://status.ovh.es/?do=details&id=6307 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Task Type Maintenance Category Backend Core Status CLOSED Percent Complete Details Maintenance works are in progress on the fibres optics of the Lille loop sont en cours sur les fibres optiques de la boucle de Lille The RSA59 and STM59 sites are currently unreachable The works should end by 6 00 Date Tuesday 11 February 2014 09 56AM Reason for closing Done Comments 1 Comment by OVH Tuesday 11 February

    Original URL path: http://status.ovh.es/?do=details&id=6295 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Task Type Modernization Category Backend Core Status CLOSED Percent Complete Details Loss of service is expected during the update reboot of the LNS 1 PAR router For more information please see https status ovh net do details id 6263 Date

    Original URL path: http://status.ovh.es/?do=details&id=6264 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    impact for SFR FT PPP customers as they were migrated to the backup LNS on Monday 03 02 2014 Date Friday 07 February 2014 09 38AM Reason for closing Done Comments 6 Comment by OVH Friday 07 February 2014 09 36AM Upgrading bootrom reboot Comment by OVH Friday 07 February 2014 09 36AM Bootrom up to date waiting for redundancy then upgrading release image Comment by OVH Friday 07 February

    Original URL path: http://status.ovh.es/?do=details&id=6263 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    11 44AM We have changed the main LNS configuration LNS for lns 1 par customers to lns 1 lyo with lns 1 rbx as a backup so that the sessions don t come back onto the Paris LNS which is causing the instabilities The PPP sessions will thus be the Lyon LNS at the time of the next reset Comment by OVH Friday 31 January 2014 11 40AM We carried

    Original URL path: http://status.ovh.es/?do=details&id=6205 (2015-10-15)
    Open archived version from archive